OpenStudy (anonymous):

Help Please! Three people are selected at random from six females and eight males. Find the probability of the following At most two are male.

OpenStudy (anonymous):

2/14= 1/7 (Not sure though)

OpenStudy (anonymous):

Or its 2/8= 1/4 ( its either of these two)

OpenStudy (anonymous):

I mean my first post and my second post

OpenStudy (uglyturtle):

I would say 1/7. I may be wrong though.

OpenStudy (anonymous):

I tried the first response earlier but I kept getting it wrong. Rounded it is approx. .143. I tried right now the 2/8 and it still marks it wrong

OpenStudy (anonymous):

U want it in decimals? I mean i have learned this all only in fractions

OpenStudy (anonymous):

Yes decimals

OpenStudy (anonymous):

Oh ok i will give it a try

OpenStudy (anonymous):

I did 2/8= 1/4 = 0.25

OpenStudy (anonymous):

Do u think its a correct answer?

OpenStudy (anonymous):

what do you thick at most 2 men means?

OpenStudy (anonymous):

No, I tried both ways and I keep getting it wrong. Thanks for your help though. =)

OpenStudy (anonymous):

you did the problem theoretically wrong, that's why the answer is wrong.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

so what does at most 2 male means?